OpenStudy (anonymous):

fan & medal Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (goformit100):

Take Matrics in use.

OpenStudy (texaschic101):

2x + y = 1 ----> y = -2x + 1 now sub -2x + 1 in for y in the other equation 4x + 2y = -1 4x + 2(-2x + 1) = -1 -- distribute 4x - 4x + 2 = -1 -- subtract 2 from both sides 4x - 4x = -1 - 2 -- combine like terms 0 = -3 (incorrect) There is no x value.There is no solution.

OpenStudy (anonymous):

thanks highly appreciated @texaschic101

OpenStudy (texaschic101):

no problem...glad to help :)